What is the total charge on an infinite wire carring a charge density


λ =


1


|x|+α where α is a cosntant.

1
Expert's answer
2021-11-08T09:12:53-0500

Given,

Charge density of the charged wire "(\lambda) = \frac{1}{|x|+\alpha}"

"\Rightarrow dq=\lambda dx"

"\Rightarrow \int dq = \int \lambda dx"

Now, let the total length of the wire be l. So, integration limit will be "-l" to "+l"

"\Rightarrow \int_{q=0}^{q=Q} dq=\int_{-l}^{l}\lambda dx"

Now, taking the integration

"Q=\int_{x=0}^{x=l} \frac{dx}{x-\alpha}-\int_{x=0}^{x=l} \frac{dx}{x+\alpha}"


"\Rightarrow Q = [\ln(x-\alpha)]_{0}^{l}-\ln{(x+\alpha)}]_{0}^{l}"


"\Rightarrow Q = \ln(x-\alpha)-ln(x+\alpha)"


"\Rightarrow Q = \ln{\frac{(x-\alpha)}{(x+\alpha)}}"


"\Rightarrow Q = \ln(\frac {x(1-\frac{\alpha}{x})}{x(1+\frac{\alpha}{x})})"

"x\rightarrow \infty"

"\Rightarrow Q = \ln 1\sim 0"

Hence, net charge on the straight infinite wire be zero.